if 4x^2-12x k is a perfect square. Find the numerical value of k.

Grade:11

1 Answers

Harshit Panwar
6 years ago
4x2-12x+k
(a-b)2=a2-2ab+b2
4x2-12x+k=a2-2ab+b2
a=2x
-2ab=-12x
b=3
therefore, k=9 for the eq to be perfect sq
Hope it helps u.